# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 2.1. Why bidimensionality fails and differences with classes of “fat” objects
Even if our goal is to abstract from a specific graph class, let us consider in this section the class of $2$-DIR graphs, corresponding to the intersection graphs of vertical or horizontal segments in the plane. As these objects are non “fat” [^5] A regions $R$ of the plane is said to be $\alpha$-fat if the radius of smallest disk enclosing $R$ is at most $\alpha$ times larger than the radius of the largest disk enclosed in $S$. A family of regions of the plane is then said to be fat if there exists $\alpha$ such that all the elements of the family are $\alpha$-fat. and can cross (unlike pseudo-disks), this class constitute a good candidate to exemplify the difficulty.
@@
A common approach is as follows. Given an instance $(G,k)$ of FVS, we compute first in polynomial time a $2$-approximation, implying that we either detect a no-instance, or define a set $M$ with $|M|\leq 2k$ and such that $G-M$ is a forest. The goal is now to reduce, using kernelization or subexponential branching rules, to an equivalent instance $(G^{\prime},k^{\prime})$ with small treewidth $\operatorname{\textsf{tw}}(G^{\prime})=k^{1-\varepsilon}$. As FVS can be solved in $\operatorname{\textsf{tw}}(G^{\prime})^{\operatorname{\textsf{tw}}(G^{\prime})}n^{\mathcal{O}(1)}$ using a classical dynamic programming approach, we get a subexponential parameterized algorithm. Thus, one has to find a way to destroy in $G$ the obstructions preventing a small treewidth. A first type of obstructions is $K_{r}$ and $K_{r,r}$, which are easy to handle as there are folklore subexponential branchings when $r=k^{\varepsilon}$. Now, one can see the hard part is destroying $K_{r,r}$ hidden (as a minor for example) (see Figure 1).
@@
Figure 1. Example of a $K_{r,r}$ contained as a minor for $r=4$ in a disk graph (left) and a $2$-DIR graph (right). In the case of disk graph, $v$ has a matching of size $r-2$ in its neighborhood, forming a triangle bundle, which can be exploited to branch. The set $M$ are depicted in blue. For the $2$-DIR graph, the vertices of the long paths are represented by segments with small variation in their height and not intersecting for better clarity, but are in fact on the same level and intersecting.
@@
A point that seems crucial to us is the following. In intersection graphs of “fat objects” (like disks, squares, or pseudo-disks more generally), the “locally non planar structure” when an object (vertex $v$ in Figure 1) is “traversed” (by $v_{1}$, $v_{2}$ in Figure 1) comes to the price of an edge ($\{v_{1},v_{2}\}$) in the neighborhood of $v$. Thus, the presence of a large $K_{r,r}$ as a minor implies that a large matching $E_{v}$ (of size $\Omega(r)$) will appear in the neighborhood of a vertex $v$. However, as $G[\{v\}\cup E_{v}]$ (called a triangle bundle in [LPS+23]) contains $r$ triangles identified on vertex $v$, the set $\{v\}\cup E_{v}$ is a good structure to perform a subexponential branching for FVS. Indeed, [LPS+22] proposed a “virtual branching” to handle this structure by either taking $v$ in the solution, or absorbing $E_{v}$ in $M$, implying then that the parameter virtually decreases by $|E_{v}|$ as a solution which does not contain $v$ has to hit all these edges, even if we cannot branch to determine which are exactly the vertices in the solution.
@@
Once no more virtual branching is possible on large triangle bundles, they obtain by some additional specialized techniques that any vertex in $M$ is such that $N_{V(G)\setminus M}(v)$ is an independent set. Then, it is proved ( [LPS+22], Corollary 1.1) that in a disk graph where for any $v\in M$, $N_{V(G)\setminus M}(v)$ is an independent set, and where there does not exist a vertex in $V(G)\setminus M$ whose neighborhood is contained in $M$, then $\operatorname{\textsf{tw}}(G)=\mathcal{O}\left(\sqrt{|M|}\omega(G)^{2.5}\right)$, where $\omega(G)$ denotes the maximum size of a clique in $G$. This no longer holds for $2$-DIR: the family of pairs $(G,M)$ depicted on the right of Figure 1 is indeed a counter example as they respect the conditions, have $\omega(G)=2$, but $\operatorname{\textsf{tw}}(G)=\Omega(|M|)$.
@@
More generally, the role of the size of a matching in the neighborhood was studied in [BBGR24a] which shows how subexponential parameterized algorithms can be obtained for graph classes having the “almost square grid minor property” (ASQGM), corresponding informally [^6] In the correct definition $\mu_{N}(G)$ is replaced by a slightly more technical parameter. to $\operatorname{\textsf{tw}}(G)=\mathcal{O}(\omega(G)\mu_{N}(G)\operatorname{\boxplus}(G))$ where $\mu_{N}(G)$ is the maximum size of a matching in a neighborhood of a vertex, and $\operatorname{\boxplus}(G)$ is the largest size of a grid contained as a minor in $G$. The previous counter example shows that $2$-DIR does not have the ASQGM property, implying that we need another approach to handle them.
@@ 2.2. A simpler case study: when trees are only paths
[…]
Let us now explain how property of item 2 (bounded tree neighborhood complexity, abbreviated bounded T-NC) allows to obtain the following “degree-related size property”: for any subpath $P$ of a $P_{i}$, $|P|\leq\operatorname{\textsf{poly}}(r)(d_{M}(P)^{\alpha+1})$. Define an independent set $\mathcal{T}$ of size $|P|/2$ by picking every second vertex in $P$. According to the definition of bounded T-NC (item 1) with $A=N_{M}(P)$, we get |$|\{N_{M}(v),~{}v\in\mathcal{T}\}|\leq f_{1}(r)(d_{M}(P))^{\alpha}$. Thus, if $|P|\geq x\cdot f_{1}(r)(d_{M}(P))^{\alpha}$, we found $x$ vertices in $P$ having the same neighborhood $M^{\prime}$ in $M$, and thus a vertex $u\in M^{\prime}$ adjacent to $x$ vertices in $P$. If $x$ is large enough (about $d_{M}(P)$), it is always better to take any arbitrary vertex $u\in M^{\prime}$. This leads to a kernelization rule (corresponding to (KR5)): if there is a vertex $u\in M$ adjacent to approximately $d_{M}(P)$ vertices in a subpath $P$, take $u$ and decrease $k$ by one. To sum it up, after applying this rule, for any subpath $P$ of a $P_{i}$, if we have $d_{M}(P)=\operatorname{\textsf{poly}}(r)$, then $|P|=\operatorname{\textsf{poly}}(r)$.
@@
Before the next step we need to apply the following “large degree rule” (corresponding to (KR3)): if there is a vertex $v$ in a $P_{i}$ such that $d_{M}(v)>t$, then add $v$ to $M$. One can prove that by taking $t=2d_{r}$, with $d_{r}=\operatorname{\textsf{poly}}(r)$ the constant defined in item 4 of the definition of a nice class, $M$ does not grow too much after applying this rule exhaustively: by denoting $A\subseteq M$ the set of vertices already in $M$ previously added by this rule, we always have $|A|=\operatorname{\textsf{poly}}(r)|M\setminus A|$. This claim will be discussed in Subsection 2.3.
@@
Observe that at this stage we may still have large $K_{r,r}$ as minor, with for example graphs as in Figure 1 where no rule applies. It remains to define a crucial rule to destroy these $K_{r,r}$. Let us now present the analogue of the Partitionning-Algorithm  of Subsection 3.7 that partitions the $P_{i}^{\prime}s$ as follows. For any connected component $P_{i}$ in $G-M$, we start (see Figure 2) from an endpoint of $P_{i}$ and collect greedily vertices until we find a subpath $P_{i}^{1}$ such that $d_{M}(P_{i}^{1})\geq t$, or that there is no more vertices in $P_{i}$. If $d_{M}(P_{i}^{1})\geq t$, then restart a new path starting from the next vertex to create $P_{i}^{2}$, and so on. This defines a partition $P_{i}=\bigcup_{\ell\in[x(P_{i})]}(P_{i}^{\ell})$, where $d_{M}(P_{i}^{\ell})\geq t$ for any $\ell\in[1,x(P_{i})-1]$ and no lower bound for $d_{M}\left(P_{i}^{x(P_{i})}\right)$. As we applied the large degree rule, we also know that $d_{M}(P_{i}^{\ell})\leq 2t$ for any $\ell\in[1,x(P_{i})]$, because collecting at each step a new vertex in the path $P_{i}^{\ell}$ can increase $d_{M}(P_{i}^{\ell})$ by at most $t$. This implies, using the degree-related size property introduced above, that $|P_{i}^{\ell}|\leq\operatorname{\textsf{poly}}(r)$ for any $\ell\in[1,x(P_{i})]$. Let us denote $\mathcal{T}^{+}$ the set of $P_{i}^{\ell}$ such that $d_{M}(P_{i}^{\ell})\geq t$ the “large-degree subpaths”. Observe that the last considered subpath $P_{i}^{x(P_{i})}$ of each connected component $P_{i}$ (on the right of each path of $G-M$ in Figure 2) may have $d_{M}\left(P_{i}^{x(P_{i})}\right)<t$ as there was no more vertices to complete it, hence it is not contained in $\mathcal{T}^{+}$. We denote $\mathcal{T}^{-}$ those remaining “small-degree subpaths”.
@@
Let us now explain how the bounded T-NC property (item 2) allows to obtain the following “small number of large-degree subpaths” property. By removing half of the $\mathcal{T}^{+}$’s, we can get a set $\mathcal{T}^{+^{\prime}}$ of non-adjacent trees (meaning with no edge between the $T_{i}^{\prime}s$) such that $|\mathcal{T}^{+^{\prime}}|\geq\frac{1}{2}|\mathcal{T}^{+}|$. We can then apply the T-NC property with $A=M$, $\mathcal{T}=\mathcal{T}^{+^{\prime}}$ and $p=m=\operatorname{\textsf{poly}}(r)$ to obtain |$|\{N_{M}(T),~{}T\in\mathcal{T}^{+^{\prime}}\}|\leq\operatorname{\textsf{poly}}(r)|M|$. Thus, if $|\mathcal{T}^{+^{\prime}}|\geq x\cdot\operatorname{\textsf{poly}}(r)|M|$, we found $x$ large-degree subpaths (denoted $T^{\prime}_{i}$) having the same neighborhood $X^{\prime}$ in $M$ with $|X^{\prime}|\geq t$. By choosing $x=t$ and considering $X\subseteq X^{\prime}$ with $|X|=t$, we found a structure that we call a $\widetilde{K_{t,t}}$ (see Figure 3), formally defined as a pair $\left(X,(T_{i})_{1\leq i\leq t}\right)$ where
@@
  • $(T_{i})_{i}$ a family of $t$ vertex-disjoint non-adjacent subtrees (paths here) of $G-M$ such that for all $1\leq i\leq t$, $X\subseteq N_{M}(T_{i})$, and
@@
Now, inspired by the “virtual branching rule” for a triangle bundle, we introduce a branching rule (corresponding to (BR2)) that either takes almost all vertices in $X$, or absorbs by adding to $M$ a subset of $t-1$ of the $T_{i}^{\prime}s$. The complexity behind this rule is fine, as in the second branch, the parameter virtually decreases by $t-1$, and $M$ grows by $(t-1)\max_{i}\{|T_{i}|\}=\operatorname{\textsf{poly}}(r)$. This explains how we deal with $K_{t,t}$ hidden as a minor.
Finally, if this $\widetilde{K_{t,t}}$ rule cannot be applied, it remains to bound $|V(G)\setminus M|$. Recall that any $P\in\mathcal{T}^{+}\cup\mathcal{T}^{-}$ is such that $|P|\leq\operatorname{\textsf{poly}}(r)$, and thus we only need to bound $|\mathcal{T}^{+}\cup\mathcal{T}^{-}|$. As we cannot apply the previous rule, we know that the number of big paths is small: $|\mathcal{T}^{+}|\leq\operatorname{\textsf{poly}}(r)|M|$. Now, to bound $|\mathcal{T}^{-}|$, observe that we can partition $\mathcal{T}^{-}=\mathcal{T}^{-}_{1}\cup\mathcal{T}^{-}_{2}$, where
@@
  • $\mathcal{T}^{-}_{2}$ is the set of small-degree path which is an entire connected component of $G-M$.
As $|\mathcal{T}^{-}_{1}|\leq|\mathcal{T}^{+}|$, it only remains to bound $|\mathcal{T}^{-}_{2}|$. Now, we can exploit once again the bounded T-NC property (item 2) to obtain that, if $|\mathcal{T}^{-}_{2}|\geq x\cdot\operatorname{\textsf{poly}}(r)|M|$, then we can find $x$ disjoint non-adjacent paths in $\mathcal{T}^{-}_{2}$ having the same neighborhood $X$ in $M$. This case is different from the $\widetilde{K_{t,t}}$ case as $X$ may be arbitrarily small (we only know that $|X|<t$), and thus unlike $\widetilde{K_{t,t}}$ this does not allow to decrease the parameter by a large amount. However, in this case, paths of $\mathcal{T}^{-}_{2}$ are just connected components, and this help us to add a last rule ((KR4)) that identifies a “redundant” path that can be safely removed. It can be shown that such a redundant path can be found just by taking $x=t+2$. Hence after applying the rule exhaustively we can assume $|\mathcal{T}^{-}_{2}|<x\cdot\operatorname{\textsf{poly}}(r)|M|=\operatorname{\textsf{poly}}(r)|M|$.
This concludes the proof for this restricted setting where the connected components of $G-M$ are paths, as we obtain by taking $\varepsilon$ small enough $|V(G)\setminus M|\leq\operatorname{\textsf{poly}}(r)|M|\leq\operatorname{\textsf{poly}}(k^{\varepsilon})k=\mathcal{O}(k^{2-\varepsilon^{\prime}})$ as required.

### v2

@@ 2.1. Why bidimensionality fails and differences with classes of “fat” objects
Even if our goal is to abstract from a specific graph class, let us consider in this overview the class of $2$-DIR graphs, corresponding to the intersection graphs of vertical or horizontal segments in the plane. As these objects are non “fat” [^5] A regions $R$ of the plane is said to be $\alpha$-fat if the radius of smallest disk enclosing $R$ is at most $\alpha$ times larger than the radius of the largest disk enclosed in $S$. A family of regions of the plane is then said to be fat if there exists $\alpha$ such that all the elements of the family are $\alpha$-fat. and can cross (unlike pseudo-disks), this class constitutes a good candidate to exemplify the difficulties.
@@
A commonly used approach is as follows. Given an instance $(G,k)$ of FVS, we compute first in polynomial time a $2$-approximation, implying that we either detect a No-instance, or define a set $M$ with $|M|\leq 2k$ and such that GMG-M is a forest. The goal is then to reduce the input, using kernelization or subexponential branching rules, to equivalent instances $(G^{\prime},k^{\prime})$ with small treewidth, that is $\operatorname{\textsf{tw}}(G^{\prime})=\mathcal{O}(k^{1-\varepsilon})$. As FVS can be solved in $2^{\operatorname{\tilde{\mathcal{O}}}(\operatorname{\textsf{tw}}(G^{\prime}))}n^{\mathcal{O}(1)}$ using a classical dynamic programming approach, we get a subexponential parameterized algorithm. Thus, one has to find a way to destroy in $G$ the obstructions preventing a small treewidth. A first type of obstructions is $K_{r}$ and $K_{r,r}$, which are easy to handle as there are folklore subexponential branchings when $r=k^{\varepsilon}$. Now, one can see the hard part is destroying a $K_{r,r}$ hidden (as a minor for example) (see Figure 1).
@@
Figure 1. Example of a $K_{r,r}$ contained as a minor for $r=4$ in a disk graph (left) and a 22-DIR graph (right). In the case of disk graph, $v$ has a matching of size r2r-2 in its neighborhood, forming a triangle bundle, which can be exploited to branch. The set $M$ are depicted in blue. For the 22-DIR graph, the vertices of the long paths are represented by segments with small variation in their height and not intersecting for better clarity, but are in fact on the same level and intersecting.
@@
A point that seems crucial to us is the following. In intersection graphs of “fat objects” (like disks, squares, or pseudo-disks more generally), the “locally non planar structure” when an object (vertex $v$ in Figure 1) is “traversed” (by $v_{1}$, $v_{2}$ in Figure 1) comes to the price of an edge ($\{v_{1},v_{2}\}$) in the neighborhood of $v$. Thus, the presence of a large $K_{r,r}$ as a minor implies that a large matching $E_{v}$ (of size $\Omega(r)$) will appear in the neighborhood of a vertex $v$. However, as $G[\{v\}\cup E_{v}]$ (called a triangle bundle in [Lokshtanov23Approx]) contains $r$ triangles pairwise intersecting on exactly one vertex $v$, the set $\{v\}\cup E_{v}$ is a good structure to perform a subexponential branching for FVS. Indeed, [lokSODA22] proposed a “virtual branching” to handle this structure by either taking $v$ in the solution, or absorbing $E_{v}$ in $M$, implying then that the parameter virtually decreases by $|E_{v}|$ as a solution which does not contain $v$ has to hit all these edges, even if we cannot branch to determine which are exactly the vertices in the solution.
@@
Once no more virtual branching is possible on large triangle bundles, they obtain by some additional specialized techniques that any vertex in $M$ is such that $N_{V(G-M)}(v)$ is an independent set. Then, it is proved ( [lokSODA22], Corollary 1.1) that in a disk graph where for any $v\in M$, $N_{V(G-M)}(v)$ is an independent set, and where there does not exist a vertex in V(GM)V(G-M) whose neighborhood is contained in $M$, then $\operatorname{\textsf{tw}}(G)=\mathcal{O}\left(\sqrt{|M|}\omega(G)^{2.5}\right)$, where $\omega(G)$ denotes the maximum size of a clique in $G$. This no longer holds for $2$-DIR graphs: the family of pairs $(G,M)$ depicted on the right of Figure 1 is indeed a counter example as they respect the conditions, have $\omega(G)=2$, but $\operatorname{\textsf{tw}}(G)=\Omega(|M|)$.
@@
More generally, the role of the size of a matching in the neighborhood was studied in [berthe24ASQGM] which shows how subexponential parameterized algorithms can be obtained for graph classes having the “almost square grid minor property” (ASQGM), corresponding more or less [^6] In the correct definition $\mu_{N}(G)$ is replaced by a slightly more technical parameter. to $\operatorname{\textsf{tw}}(G)=\mathcal{O}(\omega(G)^{\mathcal{O}(1)}\mu_{N}(G)\operatorname{\boxplus}(G))$ where $\mu_{N}(G)$ is the maximum size of a matching in a neighborhood of a vertex, and $\operatorname{\boxplus}(G)$ is the largest size of a grid contained as a minor in $G$. The previous counterexample shows that $2$-DIR does not have the ASQGM property, implying that we need another approach to handle them.
@@ 2.2. A simpler case study: when trees are only paths
To simplify the arguments, but still understand why properties in 1.5 of a nice graph class are needed, let us assume that the forest GMG-M only contains paths $(P_{i})_{i}$, and consider the example of segment graphs. This case remains challenging as a large $K_{r,r}$ can still be hidden as a minor (as in Figure 1), and we need to destroy it in order to reduce the treewidth of the graph. To keep notations simple, we use the notation $\operatorname{\textsf{poly}}(.)$ to denote a polynomial dependency on the parameter, and thus we do not try to compute tight formulas depending on the polynomial $f_{1},f_{2},f,d$ given in the definition of a nice class. We assume that we performed folklore branching and that we are left with a $K_{r,r}$-free graph $G$ for $r=k^{\varepsilon}$. It is known for string graphs (and thus segment graphs) that in this case $\operatorname{\textsf{tw}}(G)=\operatorname{\textsf{poly}}(k^{\varepsilon})n^{1/2}$ (corresponding to Item C3 of the definition of nice). Thus, our goal is to reduce $|V(G)|$ to $\mathcal{O}(k^{2-\varepsilon^{\prime}})$ for some $\varepsilon^{\prime}$. A first obvious rule is to iteratively contract edges of the $P_{i}$’s whose endpoints have no neighbors in $M$ (corresponding hereafter to (KR2)). This explains the property of Item C1 of the definition of nice. Note that after if this rule does not apply, at least half of the vertices of the $P_{i}$’s have neighbors in $M$.
Let us now explain how this rule and another rule, lead to the following “degree-related size property”: for any subpath $P$ of a $P_{i}$, $|P|=\mathcal{O}(d_{M}(P)^{2})$. Consider the subset of vertices of $P$ having neighbors in $M$, and note that it has size at least roughly $|P|/2$. This implies that there are at least $|P|/2$ edges between $P$ and $N_{M}(P)$. We now need a new kernelization rule (corresponding to (KR5)): If there exists a vertex $u\in M$ adjacent to $x$ vertices in $P$, for $x\approx 2d_{M}(P)$, then it is optimal to force $u$ in the solution. After applying this rule, for any subpath $P$ of a $P_{i}$, every vertex of $N_{M}(P)$ is adjacent to at most $2d_{M}(P)$ vertices in $P$, hence there are at most $2d_{M}(P)^{2}$ edges between $N_{M}(P)$ and $P$. Thus $|P|/2\leq 2d_{M}(P)^{2}$.
@@
Before the next step we need to apply the following “large degree rule” (corresponding to (KR3)): if there is a vertex $v$ in a $P_{i}$ such that $d_{M}(v)>t$, then add $v$ to $M$. One can prove that by taking $t=2d(r)$, with $d(r)=\operatorname{\textsf{poly}}(r)$ the constant defined in Item C4 of the definition of a nice class, $M$ does not grow too much after applying this rule exhaustively: by denoting $A\subseteq M$ the set of vertices already in $M$ before applying this rule, we always have $|A|=\operatorname{\textsf{poly}}(r)|M\setminus A|$. This claim will be discussed in Subsection 2.3.
@@
Observe that at this stage we may still have a large $K_{r,r}$ as a minor, with for example graphs as in Figure 1 where no rule applies. It remains to define a crucial rule to destroy these $K_{r,r}$-minors. Let us now present an algorithm that partitions the $P_{i}^{\prime}s$, whose analogue in the general case is called Partitionning-Algorithm  in Subsection 3.7. For any connected component $P_{i}$ in GMG-M, we start (see Figure 2) from an endpoint of $P_{i}$ and collect greedily vertices until we find a subpath $P_{i}^{1}$ such that $d_{M}(P_{i}^{1})\geq t$, or that there is no more vertices in $P_{i}$. If $d_{M}(P_{i}^{1})\geq t$, then restart a new path starting from the next vertex to create $P_{i}^{2}$, and so on. Call $x(P_{i})$ the number of subpaths defined this way from $P_{i}$. This defines a partition $P_{i}=\bigcup_{1\leq\ell\leq x(P_{i})}(P_{i}^{\ell})$, where $d_{M}(P_{i}^{\ell})\geq t$ for any $\ell\in[1,x(P_{i})-1]$ and no lower bound for $d_{M}\left(P_{i}^{x(P_{i})}\right)$. As we applied the large degree rule, we also know that $d_{M}(P_{i}^{\ell})\leq 2t$ for any $\ell\in[1,x(P_{i})]$, because collecting at each step a new vertex in the path $P_{i}^{\ell}$ can increase $d_{M}(P_{i}^{\ell})$ by at most $t$. This implies, using the degree-related size property introduced above, that $|P_{i}^{\ell}|=\mathcal{O}(t^{2})\leq\operatorname{\textsf{poly}}(r)$ for any $\ell\in[1,x(P_{i})]$. Let us denote $\mathcal{T}^{+}$ the set of those $P_{i}^{\ell}$’s such that $d_{M}(P_{i}^{\ell})\geq t$, i.e., the “large-degree subpaths”. Observe that the last considered subpath $P_{i}^{x(P_{i})}$ of each connected component $P_{i}$ (on the right of each path of GMG-M in Figure 2) may have $d_{M}\left(P_{i}^{x(P_{i})}\right)<t$ as there was no more vertices to complete it, hence it is not contained in $\mathcal{T}^{+}$. We denote $\mathcal{T}^{-}$ those remaining “small-degree subpaths”.
@@
Let us now explain how the bounded tree neighborhood complexity property (Item 2) allows to obtain the following “small number of large-degree subpaths” property. By removing half of the $\mathcal{T}^{+}$’s, we can get a set $\mathcal{T}^{+^{\prime}}$ of non-adjacent trees (meaning with no edge between the $T_{i}^{\prime}s$) such that $|\mathcal{T}^{+^{\prime}}|\geq\frac{1}{2}|\mathcal{T}^{+}|$. We can then apply the bounded tree neighborhood complexity property with $A=M$, $\mathcal{T}=\mathcal{T}^{+^{\prime}}$ and $p=m=\operatorname{\textsf{poly}}(r)$ to obtain $|\{N_{M}(T),~T\in\mathcal{T}^{+^{\prime}}\}|\leq\operatorname{\textsf{poly}}(r)|M|$. Thus, if $|\mathcal{T}^{+^{\prime}}|\geq x\cdot\operatorname{\textsf{poly}}(r)|M|$, we found $x$ large-degree subpaths (denoted $T^{\prime}_{i}$) having the same neighborhood $X^{\prime}$ in $M$ with $|X^{\prime}|\geq t$. By choosing $x=t$ and considering $X\subseteq X^{\prime}$ with $|X|=t$, we found a structure that we call a $\widetilde{K_{t,t}}$ (see Figure 3), formally defined as a pair $\left(X,\{T_{i}\}_{1\leq i\leq t}\right)$ where
@@
  • $\{T_{i}\}_{i}$ a family of $t$ vertex-disjoint non-adjacent subtrees (paths here) of GMG-M such that for all $1\leq i\leq t$, $X\subseteq N_{M}(T_{i})$, and
@@
Now, inspired by the aforementioned “virtual branching rule” of [Lokshtanov23Approx] for triangles bundles, we introduce a branching rule (corresponding to (BR2) hereafter) that either deletes almost all vertices in $X$, or is adding to $M$ a subset of t1t-1 of the $T_{i}^{\prime}s$. The complexity behind this rule is fine, as in the second branch, the parameter virtually decreases by t1t-1, and $M$ grows by $(t-1)\max_{i}\{|T_{i}|\}=\operatorname{\textsf{poly}}(r)$. This explains how we deal with the $K_{t,t}$-minors.
Finally, if this $\widetilde{K_{t,t}}$ rule cannot be applied, it remains to bound |V(GM)||V(G-M)|. Recall that any $P\in\mathcal{T}^{+}\cup\mathcal{T}^{-}$ is such that $|P|\leq\operatorname{\textsf{poly}}(r)$, and thus we only need to bound $|\mathcal{T}^{+}\cup\mathcal{T}^{-}|$. As we cannot apply the previous rule, we know that the number of big paths is small: $|\mathcal{T}^{+}|\leq\operatorname{\textsf{poly}}(r)|M|$. Now, to bound $|\mathcal{T}^{-}|$, observe that we can partition $\mathcal{T}^{-}=\mathcal{T}^{-}_{1}\cup\mathcal{T}^{-}_{2}$, where
@@
  • $\mathcal{T}^{-}_{2}$ is the set of small-degree paths which are entire connected components of GMG-M.
As $|\mathcal{T}^{-}_{1}|\leq|\mathcal{T}^{+}|$, it only remains to bound $|\mathcal{T}^{-}_{2}|$.
Let us now explain how to obtain $|\mathcal{T}^{-}_{2}|\leq\operatorname{\textsf{poly}}(r)|M|$. According to Item 2 of the bounded tree neighborhood complexity property with $\mathcal{T}=\mathcal{T}^{-}_{2}$, $A=M$, $p=t$, and $m=\operatorname{\textsf{poly}}(r)$, we get that $|\{N_{M}(T),T\in\mathcal{T}^{-}_{2}\}|\leq f_{2}(r,p,m)|M|$. Thus, by pigeonhole principle, if $|\mathcal{T}^{-}_{2}|\geq(t+2)f_{2}(r,p,m)|M|$, then there exists $t+2$ paths of $\mathcal{T}^{-}_{2}$ having the same neighborhood $X$ in $M$. This case corresponds to a $\widetilde{K_{t_{1},t_{2}}}$ where $t_{1}=|X|$ and $t_{2}=t+2$. However, as $t_{1}$ may be arbitrarily small, we cannot branch as we did on $\widetilde{K_{t,t}}$ as the branches where we delete almost all vertices of $X$ do not decrease the parameter $k$ by a large amount. However, in this case, paths of $\mathcal{T}^{-}_{2}$ are just connected components of GMG-M, and this helps to obtain a last rule ((KR4)) that identifies paths that can be safely removed. This rule, restated in our context, says that if $|X|+2$ paths in $G\setminus M$ have the same neighborhood $X$ in $M$, then one of them can be removed. As in our case we found $t+2$ paths having the same neighborhood $X$, and $t>|X|$, we can indeed apply the rule. Thus, once (KR4) cannot be applied, we get $|\mathcal{T}^{-}_{2}|<(t+2)f_{2}(r,p,m)|M|\leq\operatorname{\textsf{poly}}(r)|M|$.
This concludes the sketch of proof for this restricted setting where the connected components of GMG-M are paths, as we obtain by taking $\varepsilon$ small enough $|V(G-M)|\leq\operatorname{\textsf{poly}}(r)|M|\leq\operatorname{\textsf{poly}}(k^{\varepsilon})k=\mathcal{O}(k^{2-\varepsilon^{\prime}})$ as required.